# Shortcuts and Identifiability in Concept-based Models from a Neuro-Symbolic Lens

### v1

@@ 5 Case Studies
We tackle the following key research questions:
  • Q1.
    Are CBMs affected by JRSs in practice?
  • Q2.
    Do JRSs affect interpretability and OOD behavior?
  • Q3.
    Can existing mitigation strategies prevent JRSs?
Additional implementation details about the tasks, architectures, and model selection are reported in Appendix A.
Models. We evaluate several (also NeSy) CBMs. DeepProbLog (DPL[Manhaeve et al., 2018, 2021] is the only method supplied with the ground-truth knowledge $\mathsf{K}$, and uses probabilistic-logic reasoning to ensure predictions are consistent with it. CBNM is a Concept-Bottleneck Model [Koh et al., 2020] with no supervision on the concepts. Deep Symbolic Learning (DSL[Daniele et al., 2023] is a SOTA NeSy-CBM that learns concepts and symbolic knowledge jointly; it implements the latter as a truth table and reasoning using fuzzy logic. We also consider two variants: DPL replaces fuzzy with probabilistic logic, while bears wraps DPL within BEARS [Marconato et al., 2024], an ensemble approach for handling regular RSs. In short, bears consists of a learned inference layer and multiple concept extractors. The former is learned only once, the latter are learned sequentially. We also evaluate Self-explainable Neural Networks (SENN) [Alvarez-Melis and Jaakkola, 2018], unsupervised CBMs that include a reconstruction penalty. Since they do not satisfy our assumptions, they are useful to empirically assess the generality of our remarks.
Data sets. In order to assess both learned concepts and inference layer, we use three representative NeSy tasks with well-defined concept annotations and prior knowledge. MNIST-Add [Manhaeve et al., 2018] involves predicting the sum of two MNIST digits, and serves as a sanity check as it provably contains no RSs. MNIST-SumParity is similar except we have to predict the parity of the sum, as in Fig. 1, and admits many JRSs. We also include a variant of Clevr [Johnson et al., 2017] in which images belong to $3$ classes as determined by a logic formula and only contain $2$ to $4$ objects, for scalability.
Metrics. For each CBM, we evaluate predicted labels and concepts with the $F_{1}$ score (resp. $\mathrm{F_{1}}(Y)$ and $\mathrm{F_{1}}(C)$) on the test split. Computing $\mathrm{F_{1}}(C)$ requires to first align the predicted concepts to ground-truth annotations. We do so by using the Hungarian algorithm [Kuhn, 1955] to find a permutation $\pi$ that maximizes the Pearson correlation across concepts. Concept collapse $\mathsf{Cls}(C)$ quantifies to what extent the concept extractor blends different ground-truth concepts together: high collapse indicates that it predicts fewer concepts than expected. We measure the quality of the learned inference layer $\bm{\mathrm{\beta}}$ as follows: for each example, we reorder the concept annotations $\bm{\mathrm{g}}$ using $\pi$, feed the result to $\bm{\mathrm{\beta}}$ and compute the $F_{1}$ score ($\mathrm{F_{1}}(\bm{\mathrm{\beta}})$) of its predictions. See Section A.4 for details.
Mitigation strategies. We evaluate all strategies discussed in Section 4 as well as a combination of all unsupervised mitigations, denoted H+R+C (entropy, reconstruction, contrastive). Implementations are taken verbatim from [Bortolotti et al., 2024]. For contrastive learning, we apply an InfoNCE loss to the predicted concept distribution with the augmentations suggested by Chen et al. [2020a]. For knowledge distillation, the inference layer is fed ground-truth concepts and trained to optimize the log-likelihood of the corresponding labels.
@@
Method $\mathrm{F_{1}}(Y)$ ($\uparrow$) $\mathrm{F_{1}}(C)$ ($\uparrow$) $\mathsf{Cls}(C)$ ($\downarrow$) $\mathrm{F_{1}}(\bm{\mathrm{\beta}})$ ($\uparrow$)
DPL $0.98\pm 0.01$ $0.99\pm 0.01$ $0.01\pm 0.01$-
CBNM $0.98\pm 0.01$ $0.99\pm 0.01$ $0.01\pm 0.01$ $1.00\pm 0.01$
SENN $0.97\pm 0.01$ $0.80\pm 0.07$ $0.01\pm 0.01$ $0.75\pm 0.08$
DSL $0.96\pm 0.02$ $0.98\pm 0.01$ $0.01\pm 0.01$ $1.00\pm 0.01$
DPL $0.91\pm 0.09$ $0.92\pm 0.08$ $0.01\pm 0.01$ $0.90\pm 0.08$
bears $0.76\pm 0.15$ $0.87\pm 0.13$ $0.01\pm 0.01$ $0.67\pm 0.14$
@@
Method $\mathrm{F_{1}}(Y)$ ($\uparrow$) $\mathrm{F_{1}}(C)$ ($\uparrow$) $\mathsf{Cls}(C)$ ($\downarrow$) $\mathrm{F_{1}}(\bm{\mathrm{\beta}})$ ($\uparrow$)
@@
Method $\mathrm{F_{1}}(Y)$ ($\uparrow$) $\mathrm{F_{1}}(C)$ ($\uparrow$) $\mathsf{Cls}(C)$ ($\downarrow$) $\mathrm{F_{1}}(\bm{\mathrm{\beta}})$ ($\uparrow$)
@@
C-Sup K-Sup $\mathrm{F_{1}}(Y)$ ()($\uparrow$)( ↑ ) $\mathrm{F_{1}}(\bm{\mathrm{\beta}})$ ()($\uparrow$)( ↑ ) $\mathrm{F_{1}}(Y)$ ()($\uparrow$)( ↑ ) $\mathrm{F_{1}}(\bm{\mathrm{\beta}})$ ()($\uparrow$)( ↑ )
$0\%$ $0\%$ $0.99\pm 0.01$ $0.55\pm 0.05$ $0.01\pm 0.01$ $0.47\pm 0.07$
$0\%$ $100\%$ $0.99\pm 0.01$ $0.56\pm 0.06$ $0.01\pm 0.01$ $0.58\pm 0.08$
@@
$100\%$ $100\%$ $0.97\pm 0.01$ $0.95\pm 0.01$ $0.97\pm 0.02$ $0.69\pm 0.27$
Figure 3: Traditional mitigations have limited effect on CBNM (top) and DPL (bottom) for MNIST-SumParity. The only outlier is contrastive learning (orange and purple lines), which consistently ameliorates concept collapse.
@@
Q1: All tested CBMs suffer from JRSs and simplicity bias. We analyze three learning tasks of increasing complexity. MNIST-Add provably contains no JRSs and thus satisfies the preconditions of Theorem 3.8. Compatibly with this result, all CBMs attain high label performance ($\mathrm{F_{1}}(Y)\geq 90\%$) and high-quality semantics, as shown by the values of $\mathrm{F_{1}}(C)$ ($\geq 90\%$) and $\mathrm{F_{1}}(\bm{\mathrm{\beta}})$ ($\geq 75\%$) seen in Table 1. The only exception is bears, which by design trades off performance for calibration. However, in MNIST-SumParity and Clevr, [^4] For Clevr, we focus on representative CBMs with high fit on the training data. which are more complex, all models lack intended semantics. This is clear from Tables 3 and 2: despite attaining high $\mathrm{F_{1}}(Y)$ ($\geq 90\%$), the learned concepts and inference layer are both low quality: $\mathrm{F_{1}}(C)$ is always low – it never crosses the $50\%$ mark – and $\mathrm{F_{1}}(\bm{\mathrm{\beta}})$ is at best around chance level.
@@
The behavior of concept collapse strongly hints at simplicity bias. While no collapse takes place in MNIST-Add ($\mathsf{Cls}(C)\leq 0.01$, i.e., digits are not compacted together), in MNIST-SumParity and Clevr collapse is omnipresent ($\mathsf{Cls}(C)\geq 0.22$), suggesting that CBMs are prone to simplicity bias, as expected. SENN is an outlier, likely due to the higher capacity of its inference layer (rather than reconstruction, which is entirely ineffective in Q3). Even having access to the ground-truth knowledge is ineffective, as show by the high degree of collapse displayed by DPL.
@@
Q2: JRSs compromise OOD behavior and supervision only partially helps. We evaluate the impact of JRSs and supervision on OOD behavior by training a CBNM on the biased version of MNIST-SumParity in Fig. 1. The in-distribution (ID) data comprise all (odd, odd), (even, even), and (odd, even) pairs of MNIST digits, while the OOD split contains all (even, odd) pairs. Table 4 lists results on both ID and OOD test sets for varying percentages of concept and knowledge supervision. While all models produce high-quality predictions in-distribution ($\mathrm{F_{1}}(Y)\geq 0.97$), only the CBNM receiving complete supervision attains acceptable OOD predictions: for the others, $\mathrm{F_{1}}(Y)\leq 0.40$. Even in this case, though, the inference layer still does not have intended semantics, as shown by $\mathrm{F_{1}}(\bm{\mathrm{\beta}})\leq 70\%$.
@@
Q3: Popular CBM and RS mitigations fall short. Finally, we apply each mitigation strategy to a CBNM trained on MNIST-SumParity and Clevr, ablating the amount of supervision. The results for MNIST-SumParity in Fig. 3 show while concept supervision ($x$-axis) helps all metrics, adding reconstruction, entropy maximization, and contrastive learning to the mix brings no benefits for concept ($\mathrm{F_{1}}(C)$) and knowledge ($\mathrm{F_{1}}(\bm{\mathrm{\beta}})$) quality, not even when combined (orange curve). Knowledge supervision is similarly ineffective (blue lines). This is not unexpected: MNIST-SumParity and Clevr suffer from RSs even when the model is supplied the prior knowledge. The only noticeable improvement comes from contrastive learning for concept collapse ($\mathsf{Cls}(C)$): the orange and purple curves in Fig. 3 (right) show it can effectively prevent CBMs from mixing concepts together. The results for Clevr in Appendix B paint a similar picture.
@@ 6 Related Work
Shortcuts in machine learning. Shortcuts occur when machine learning models solve a prediction task by using features that correlate with but do not cause the desired output, leading to poor OOD behavior [Geirhos et al., 2020, Teso et al., 2023, Ye et al., 2024, Steinmann et al., 2024a]. The shortcuts literature however focuses on black-box models rather than CBMs. Existing studies on the semantics of CBM concepts [Margeloiu et al., 2021, Mahinpei et al., 2021, Havasi et al., 2022, Raman et al., 2023] focus on individual failures but propose no well-defined, formal notion of intended semantics. One exception is [Marconato et al., 2023a] which however, like other works, ignores the role of the inference layer. We build on known results on reasoning shortcuts [Li et al., 2023, Marconato et al., 2023b, Wang et al., 2023, Umili et al., 2024] which are restricted to NeSy-CBMs in which the prior knowledge is given and fixed. Our analysis upgrades these results to general CBMs. Furthermore, our simulations indicate that strategies that are effective for CBMs and RSs no longer work for JRSs. At the same time, while NeSy approaches that learn the prior knowledge and concepts jointly are increasingly popular [Wang et al., 2019, Yang et al., 2020, Liu et al., 2023, Daniele et al., 2023, Tang and Ellis, 2023, Wüst et al., 2024], existing work neglects the issue of shortcuts in this setting. Our work partially fills this gap.
Relation to identifiability. Works in Independent Component Analysis and Causal Representation Learning focus on the identifiability of representations up to an equivalence relation, typically up to permutations and rescaling [Khemakhem et al., 2020a, Gresele et al., 2020, von Kügelgen et al., 2024] or more general transformations [Roeder et al., 2021, Buchholz et al., 2022]. The equivalence relation introduced along with intended semantics (Definition 3.3) establishes a specific connection between the maps $\bm{\mathrm{\alpha}}$ and $\bm{\mathrm{\beta}}$. This aligns with existing works that aim to identify representations and linear inference layers using supervised labels [Lachapelle et al., 2023, Fumero et al., 2023, Bing et al., 2023]. Taeb et al. [2022] studied the identifiability of a VAE-based CBM with continuous-valued concepts. In contrast, we provide an analysis of a broader class of (NeSy) CBMs commonly used in practice. Moreover, while prior works primarily focus on identifying continuous-valued representations, we examine the identifiability of discrete concepts and of the inference layer. [Rajendran et al., 2024] explores the identifiability of (potentially discrete) concepts linearly encoded in model representations but is limited to a multi-environment setting, not including supervised labels. Our results differ in that we show how concepts can be identified in CBMs by circumventing joint reasoning shortcuts.
@@ 7 Conclusion
We study the issue of learning CBMs and NeSy-CBMs with high-quality quality concepts and inference layers. To this end, we formalize intended semantics and joint reasoning shortcuts (JRSs), show how reducing the number deterministic JRSs to zero provably prevents all JRSs, yielding identifiability, and investigate natural mitigation strategies. Numerical simulations indicate that JRSs are very frequent and impactful, and that the only (partially) beneficial mitigation is contrastive learning. Our analysis aims to pave the way to the design of effective solutions and therefore more robust and interpretable CBMs.
In future work, we plan to extend our theory to concept learning in general neural nets and large language models. Moreover, we will take a closer look at more complex mitigations such as techniques for smartly constraining the inference layer [Liu et al., 2023, Wüst et al., 2024], debiasing [Bahadori and Heckerman, 2021], and interaction with human experts [Teso et al., 2023].

### v2

@@ 5 Case Studies
Method $F_{1}(Y)$ ($\uparrow$) $F_{1}(C)$ ($\uparrow$) $\mathsf{Cls}(C)$ ($\downarrow$) $F_{1}(\bm{\mathrm{\beta}})$ ($\uparrow$)
DPL $8\pm 0.01 0.99±0$$9\pm 0.01$ $0.01\pm 0.01$ -
[…]
SENN $0.97\pm 0.01$ $80\pm 0.07 0.0$$01\pm 0.01$ $0.75\pm 0.08$
[…]
DPL $0.91\pm 0.09$ $92\pm 0.08 0.0$$01\pm 0.01$ $0.90\pm 0.08$
[…]
@@
Method $F_{1}(Y)$ ($\uparrow$) $F_{1}(C)$ ($\uparrow$) $\mathsf{Cls}(C)$ ($\downarrow$) $F_{1}(\bm{\mathrm{\beta}})$ ($\uparrow$)
@@
Method $F_{1}(Y)$ ($\uparrow$) $F_{1}(C)$ ($\uparrow$) $\mathsf{Cls}(C)$ ($\downarrow$) $F_{1}(\bm{\mathrm{\beta}})$ ($\uparrow$)
@@
C-Sup K-Sup $F_{1}(Y)$ $(\uparrow)$ $F_{1}(\bm{\mathrm{\beta}})$ $(\uparrow)$ $F_{1}(Y)$ $(\uparrow)$ $F_{1}(\bm{\mathrm{\beta}})$ $(\uparrow)$
$\% 00\%$ $0.99\pm 0.01$ $55\pm 0.05 0.0$$01\pm 0.01$ $0.47\pm 0.07$
[…]
@@
[…]
Table 5: Results for BDD-OIA.
Method $F_{1}(Y)$ ($\uparrow$) $F_{1}(C)$ ($\uparrow$) $\mathsf{Cls}(C)$ ($\downarrow$) $F_{1}(\bm{\mathrm{\beta}})$ ($\uparrow$)
DPL $0.69\pm 0.03$ $0.44\pm 0.01$ $0.88\pm 0.01$ -
[…]
We tackle the following key research questions: Q1. Are CBMs affected by JRSs in practice? Q2. Do JRSs affect interpretability and OOD behavior? Q3. Can existing mitigation strategies prevent JRSs? Appendix˜A reports additional details about the tasks, architectures, and model selection.
Models. We evaluate several (also NeSy) CBMs. DeepProbLog (DPL) [41, 79] is the only method supplied with the ground-truth knowledge $\mathsf{K}$, and uses probabilistic-logic reasoning to ensure predictions are consistent with it. CBNM is a Concept-Bottleneck Model [3] with no supervision on the concepts. Deep Symbolic Learning (DSL) [50] is a SOTA NeSy-CBM that learns concepts and symbolic knowledge jointly; it implements the latter as a truth table and reasoning using fuzzy logic. We also consider two variants: DPL replaces fuzzy with probabilistic logic, while bears wraps DPL within BEARS [63], an ensemble approach for handling regular RSs. In short, bears consists of a learned inference layer and multiple concept extractors. The former is learned only once, the latter are learned sequentially. We also evaluate Self-explainable Neural Networks (SENN) [1], unsupervised CBMs that include a reconstruction penalty. Since they do not satisfy our assumptions, they are useful to empirically assess the generality of our remarks. In our experiments, we train all CBM variants in a joint manner [3] using only label supervision and do not rely on foundation models for gathering annotations [32, 33, 34]. For additional details, see Appendix˜A.
Data sets. To assess both learned concepts and inference layer, we use three representative NeSy tasks with explicit concept annotations and prior knowledge. MNIST-Add [41] involves predicting the sum of two MNIST digits, and serves as a sanity check as it provably contains no RSs. MNIST-SumParity is similar except we have to predict the parity of the sum, as in Fig.˜1, and admits many JRSs. We also include a variant of Clevr [83] in which images belong to $3$ classes as determined by a logic formula and only contain $2$ to $4$ objects, for scalability. Finally, we carry out basic checks also on BDD-OIA [84], a real-world autonomous driving task where images are annotated with actions (move_forward, stop, turn_left, turn_right) and 21 binary concepts.
Metrics. For each CBM, we evaluate predicted labels and concepts with the $F_{1}$ score (resp. $F_{1}(Y)$ and $F_{1}(C)$) on the test split. Computing $F_{1}(C)$ requires to first align the predicted concepts to ground-truth annotations. We do so by using the Hungarian algorithm [64] to find a permutation $\pi$ that maximizes the Pearson correlation across concepts. In this way, we directly evaluate disentanglement of the concepts as per Eq.˜5 (left). Concept collapse $\mathsf{Cls}(C)$ quantifies to what extent the concept extractor blends different ground-truth concepts together: high collapse indicates that it predicts fewer concepts than expected. We measure the quality of the learned inference layer $\bm{\mathrm{\beta}}$ as follows: for each input $\bm{\mathrm{x}}$, (1) we reorder the corresponding concept annotations $\bm{\mathrm{g}}$ using $\pi$, (2) feed the result to $\bm{\mathrm{\beta}}$, and (3) compute the $F_{1}$-score ($F_{1}(\bm{\mathrm{\beta}})$) of its predictions. Step (1) permutes and applies element-wise invertible transformations on the ground-truth concepts using the inverse of the map returned by Hungarian matching. This evaluates whether the learned inference layer can correctly handle (an invertible transformation of) the ground-truth concept annotations, measuring the generalization as per Eq.˜5 (right).This procedure is precisely detailed in Section˜A.6.
Mitigation strategies. We evaluate all strategies discussed in Section˜4 as well as a combination of all unsupervised mitigations, denoted H+R+C (entropy, reconstruction, contrastive). Implementations are taken verbatim from [22]. For contrastive learning, we apply an InfoNCE loss to the predicted concept distribution with the augmentations suggested by Chen et al. [80]. For knowledge distillation, the inference layer is fed ground-truth concepts and trained to optimize the log-likelihood of the corresponding labels.
@@
Q1: All tested CBMs suffer from JRSs and simplicity bias. We analyze three learning tasks of increasing complexity. MNIST-Add provably contains no JRSs and thus satisfies the preconditions of Theorem˜3.9. Compatibly, all CBMs attain high label performance ($F_{1}(Y)\geq 90\%$) and high-quality semantics, as shown by the values of $F_{1}(C)$ ($\geq 90\%$) and $F_{1}(\bm{\mathrm{\beta}})$ ($\geq 75\%$) seen in Table˜5. The only exception is bears, which by design trades off performance for calibration. However, in MNIST-SumParity and Clevr, [^4] For Clevr, we focus on representative CBMs with high fit on the training data. which are more complex, all models lack intended semantics. This is clear from Tables˜5 and 5: despite attaining high $F_{1}(Y)$ ($\geq 90\%$), the learned concepts and inference layer are low quality: $F_{1}(C)$ is always low – it never crosses the $50\%$ mark – and $F_{1}(\bm{\mathrm{\beta}})$ is at best around chance level. Our results on BDD-OIA in Table˜5 show the very same trend.
@@
The behavior of concept collapse strongly hints at simplicity bias. While no collapse takes place in MNIST-Add ($\mathsf{Cls}(C)\leq 0.01$, i.e., digits are not compacted together), in MNIST-SumParity and Clevr collapse is omnipresent ($\mathsf{Cls}(C)\geq 0.22$), suggesting that CBMs are prone to simplicity bias, as expected. SENN is an outlier, likely due to the higher capacity of its inference layer (rather than reconstruction, which is entirely ineffective in Q3). Even having access to the ground-truth knowledge is ineffective, as shown by the high degree of collapse displayed by DPL.
@@
Q2: JRSs compromise OOD behavior and supervision only partially helps. We evaluate the impact of JRSs and supervision on OOD behavior by training a CBNM on the biased version of MNIST-SumParity in Fig.˜1. The in-distribution (ID) data comprise all (odd, odd), (even, even), and (odd, even) pairs of MNIST digits, while the OOD split contains all (even, odd) pairs. Table˜5 reports performance under varying levels of concept and knowledge supervision. While all models produce high-quality predictions in-distribution ($F_{1}(Y)\geq 0.97$), only the CBNM receiving complete supervision attains acceptable OOD predictions: for the others, $F_{1}(Y)\leq 0.40$. Even in this case, though, the inference layer still does not have intended semantics, as shown by $F_{1}(\bm{\mathrm{\beta}})\leq 70\%$.
@@
Q3: Popular CBM and RS mitigations fall short. Finally, we test mitigation strategies on CBNM trained on MNIST-SumParity and Clevr, ablating the amount of supervision. The results for MNIST-SumParity in Fig.˜3 show while concept supervision ($x$-axis) helps all metrics, adding reconstruction, entropy maximization, and contrastive learning to the mix brings no benefits for concept and knowledge ($F_{1}(C)$, $F_{1}(\bm{\mathrm{\beta}})$) quality, not even when combined (orange curve). Knowledge supervision is also ineffective (blue lines), likely because MNIST-SumParity and Clevr suffer from RSs even when the model is supplied prior knowledge. Contrastive learning improves concept collapse ($\mathsf{Cls}(C)$): the orange and purple curves in Fig.˜3 (right) show it can prevent CBMs from mixing concepts together. The results for Clevr in Appendix˜B show similar trends.
Figure 3: Traditional mitigations have limited effect on CBNM for MNIST-SumParity. The only outlier is contrastive learning (orange and purple), which consistently ameliorates concept collapse.
@@ 6 Related Work
Shortcuts in machine learning. Shortcuts occur when machine learning models solve a prediction task by using features that correlate with but do not cause the desired output, leading to poor OOD behavior [12, 10, 85, 86]. Existing work focuses on black-box models rather than CBMs. Existing studies on the semantics of CBM concepts [87, 15, 88, 89] focus on individual failures but lack formal notion of intended semantics. One exception is [17] which, however, ignores the role of the inference layer. We build on known results on reasoning shortcuts [90, 19, 20, 53] which are restricted to NeSy-CBMs in which the prior knowledge is given and fixed. Our analysis upgrades these results to general CBMs. Furthermore, our simulations indicate that strategies that are effective for CBMs and RSs no longer work for JRSs. At the same time, while NeSy approaches that learn prior knowledge and concepts jointly are increasingly popular [48, 91, 49, 50, 51, 52], existing work neglects the issue of shortcuts in this setting. Our work partially fills this gap. Shortcuts and JRSs are related but distinct. While shortcuts can induce JRSs, the converse is not true: a model can be affected by a JRS even if it does not rely on confounders, as shown in Fig.˜2. This distinction also implies that mitigation strategies designed for vanilla shortcuts do not necessarily transfer to JRSs.
Relation to identifiability. Works in Independent Component Analysis and Causal Representation Learning focus on the identifiability of representations up to an equivalence relation, typically up to permutations and rescaling [92, 93, 69] or more general transformations [94, 95]. The equivalence relation introduced along with intended semantics (Definition˜3.3) establishes a specific connection between the maps $\bm{\mathrm{\alpha}}$ and $\bm{\mathrm{\beta}}$. This aligns with existing works that aim to identify representations and linear inference layers using supervised labels [68, 96, 97, 98, 99]. Moreover, while prior works primarily focus on identifying real-valued representations, we examine the identifiability of discrete concepts and of the inference layer. [100] explores the identifiability of (potentially discrete) concepts linearly encoded in model representations but focuses on multi-environment settings. Our results differ in that we show how concepts can be identified in CBMs by circumventing reasoning shortcuts.
@@ 7 Conclusion
We study the issue of learning CBMs and NeSy-CBMs with high-quality concepts and inference layers. We formalize intended semantics and joint reasoning shortcuts (JRSs), showing how, under suitable assumptions, zeroing the number of deterministic JRSs provably prevents all JRSs, yielding identifiability. Numerical simulations indicate that JRSs are very frequent and impactful, and that the only (partially) beneficial mitigation stratgy is contrastive learning. Our work paves the way to the design of effective solutions and therefore more robust and interpretable CBMs. In future work, we plan to extend our theory to general neural nets and large language models, and we will take a closer look at more complex mitigations such as smartly constraining the inference layer [49, 52], debiasing [13], and human interaction [10]. Another asset is considering concept-level supervision gathered from foundation models like CLIP, however recent results [35] call for caution in using them to supervise the bottleneck. Wrong concept-level supervision may transfer biases of foundation models and induce JRSs, thus resulting ineffective as a mitigation strategy.
Broader impact. With this work, we aim to highlight the issue of joint reasoning shortcuts in concept-based models and their subtle but impactful consequences on the trustworthiness and interpretability of these models, as well as on their reliability in out-of-distribution scenarios. It also highlights the limited effect of unsupervised mitigation strategies, thus pointing out a significant gap in our toolbox for addressing joint reasoning shortcuts effectively and cheaply. Our work also provides a solid theoretical foundation upon which further studies can build.
Limitations. Our theoretical analysis of JRSs builds on two common assumptions for the data generation process, which limit its scope. Relaxing them, while not straightforward, will allow treating models and NeSy tasks with intrinsic uncertainty on label and concepts. Our experiments span both synthetically-generated and a real-world vision datasets; further analysis should highlight the impact of JRSs in other real-world settings and beyond the vision domain.